OpenStudy (anonymous):

Snookers lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three time as many units of plywood as lumberr. It make a profit of $30 on a unit of lumber and $40 on a unit of plywood. How many of each units must be produced and sold in order to make a profit of $15,600? Will give a medal for the answer!!!!

OpenStudy (anonymous):

lumber = x plywood = 3x 30(x) + 40(3x) = 15600 30x + 120x = 15600 150x = 15600 x = 104 which means 104 units of lumber 3x = 3(104) = 312 which means 312 units of plywood

OpenStudy (anonymous):

thanks would you like another????